### v2

# Quantum typicality approach to energy flow between two spin-chain domains at different temperatures

Abstract
We discuss a quantum typicality approach to examine systems composed of two subsystems at different temperatures. While dynamical quantum typicality is usually used to simulate high-temperature dynamics, we also investigate low-temperature dynamics using the method. To test our method, we investigate the energy current between subsystems at different temperatures in various paradigmatic spin-1/2 chains, specifically the XX chain, the critical transverse-field Ising chain, and the XXZ chain. We compare our numerics to existing analytical results and find a convincing agreement for the energy current in the steady state for all considered models and temperatures.
I Introduction
Over the last decades, the broad field of quantum many-body systems out of equilibrium has crystallized as one of the central topics of modern physics, both experimentally and theoretically Bloch et al. [2008], Polkovnikov et al. [2011], Eisert et al. [2015], D’Alessio et al. [2016], Abanin et al. [2019], ranging from fundamental questions in statistical mechanics to applied questions in material science. The understanding of nonequilibrium physics has developed remarkably over the last years due to experimental advances Bloch et al. [2008], new theoretical concepts such as the typicality of pure states Alben et al. [1975], De Raedt and de Vries [1989], Jaklič and Prelovšek [1994], Hams and De Raedt [2000], Gemmer and Mahler [2003], Iitaka and Ebisuzaki [2003], Gemmer et al. [2004], Goldstein et al. [2006], Popescu et al. [2006], Reimann [2007], Bartsch and Gemmer [2009], Sugiura and Shimizu [2012, 2013], Elsayed and Fine [2013], Steinigeweg et al. [2014], Monnai and Sugita [2014], Endo et al. [2018], Richter and Steinigeweg [2019], Wietek et al. [2019], Rousochatzakis et al. [2019], Heitmann et al. [2020], Schnack et al. [2020], Jin et al. [2021], Reimann [2018], Mitrić [2025] and eigenstate thermalization Deutsch [1991], Srednicki [1994], Rigol et al. [2008], and the development of sophisticated numerical techniques Jaklič and Prelovšek [2000], Long et al. [2003], Schollwöck [2005], Grossjohann and Brenig [2010], Schollwöck [2011], De Raedt et al. [2017].
One paradigmatic example of a nonequilibrium process is transport Bertini et al. [2021], a natural phenomenon in systems with one or more conserved quantities. Additionally, transport appears in both open and closed systems. In an open system scenario, transport processes are usually induced by coupling the system at its boundaries to external baths at different temperatures or chemical potentials. Then, in the long-time limit, a nonequilibrium steady state usually emerges with a characteristic density profile and a constant current Michel et al. [2003], Prosen and Žnidarič [2009], Žnidarič [2011], Xu et al. [2023]. Such a situation is often modeled by a quantum master equation of Lindblad form Breuer and Petruccione [2007], which enables the application of sophisticated numerical techniques.
In a closed system scenario, a widely used approach for studying transport is linear response theory, where the Kubo formula in terms of autocorrelation functions plays a crucial role Kubo et al. [1991]. This theory predicts the behavior of a system near equilibrium, and it can be formulated either in frequency and momentum space or in the space and time domain. Another approach for studying transport in closed systems is a quantum quench Calabrese and Cardy [2006], Mitra [2018], Richter et al. [2018]. Here, a quantum system is prepared in a particular initial state, typically an eigenstate of some pre-quench Hamiltonian. Then, the system undergoes a sudden change and is described by a new Hamiltonian, and the initial state evolves unitarily according to this post-quench Hamiltonian. Quantum quenches can be studied in bipartite systems Bertini et al. [2016], Fischer et al. [2020], where the total system is composed of two subsystems and initially both subsystems are prepared in a state according to their respective Hamiltonians. Then, the subsystems are brought into contact together, and the total state evolves under the Hamiltonian of the total system.
Numerically, transport in open and closed systems has been investigated using various methods, including exact and Lanczos diagonalization Jaklič and Prelovšek [2000], Long et al. [2003], Karrasch et al. [2013], simulations based on matrix product states Vidal [2004], Zwolak and Vidal [2004], Verstraete et al. [2008], Prosen and Žnidarič [2009], Schollwöck [2011], Weimer et al. [2021], Paeckel et al. [2019], and Monte-Carlo techniques Alvarez and Gros [2002], Michel et al. [2008], Grossjohann and Brenig [2010], De Raedt et al. [2017], to name a few. Additionally, the method of dynamical quantum typicality Heitmann et al. [2020], Jin et al. [2021] is a multifaceted method for investigating transport in closed quantum systems. The basic idea of dynamical quantum typicality is that one can imitate the expectation value of an entire ensemble by an expectation value of a single pure state, drawn at random from a high-dimensional Hilbert space. Dynamical quantum typicality has been used for linear-response functions, mostly at high temperatures Bertini et al. [2021]. For such temperatures, the statistical error is smaller than at low temperatures. Still, dynamical quantum typicality is a priori neither restricted to linear-response functions nor high temperatures.
In this work, we demonstrate the usefulness of dynamical quantum typicality in a specific physical situation, which differs from most previous applications in two ways. First, we apply it to energy flow in bipartite systems. Second, we also explore low temperatures. We do so for various spin-1/2 chains, which all feature ballistic energy transport: the XX chain, the critical transverse-field Ising chain, and the XXZ chain. We compare our numerical results to existing analytical results from conformal field theory (CFT) Bernard and Doyon [2012], Fischer et al. [2020] and generalized hydrodynamics (GHD) Bertini et al. [2016], Castro-Alvaredo et al. [2016], Doyon et al. [2025]. We find convincing agreement in all considered subsystem configurations of our bipartite setup, as well as for all examined temperatures. This agreement is not only found for the contact current but also for all local currents and densities in the steady state.
Our paper is structured as follows: In Sec. II, we define the bipartite setup used throughout the paper and the energy current between the two subsystems. After that, we introduce the different spin-$1/2$ chains in Sec. III, which we use for our subsystem configurations. Next, in Sec. IV, we discuss our quantum typicality approach to the energy current and briefly review the CFT results. Thereafter, we present our numerical results and compare them to analytical results in Sec. V. We conclude in Sec. VI and give directions for further research.
II Setup
Figure 1: Schematic representation of the setup. The total system consists of two subsystems described by $H_{L}$ and $H_{R}$, respectively. Initially, they are prepared at temperatures $T_{L}$ and $T_{R}$. The interaction of both parts for times $tJ\geq 0$ is described by the Hamiltonian $H_{C}$.
Throughout our paper, we consider a system comprising two subsystems. Thus, the full Hamiltonian is written as
$H=H_{L}+H_{R}+H_{C}\;,$ (1)
where $H_{L}$ and $H_{R}$ are the Hamiltonian of the left subsystem and right subsystem, respectively. The term $H_{C}$ describes the interaction between both subsystems. By $L_{L}$ and $L_{R}$ we denote the number of sites of the two subsystems such that $L=L_{L}+L_{R}$ is the total number of sites of the system. Here, we only focus on the equally sized case of $L_{L}=L_{R}=L/2$. A schematic representation of our setup is shown in Fig. 1.
In our paper, we are interested in the heat current from the left half into the right one. This heat current is given by Giazotto et al. [2006]
$Q_{L}=J_{0}-\mu\dot{N}_{L}\;,$ (2)
where $J_{0}=-\dot{E}_{L}$ denotes the energy current flowing out of the left subsystem, $\dot{N}_{L}$ is the particle current, and $\mu$ is a chemical potential. We focus on the case of $\mu=0$. Thus, the heat current equals the energy current, $Q_{L}=J_{0}$. By using the Heisenberg equation of motion $(\hbar=1)$, we can express the energy current $J_{0}$ as
$J_{0}=-\dot{E}_{L}=-\frac{\text{d}}{\text{d}t}\langle H_{L}\rangle=i\langle[H_{L},H]\rangle\;,$ (3)
where the brackets denote the expectation value with respect to some state.
Since we are interested in the energy current $J_{0}$, we prepare both subsystems initially in a canonical ensemble with different low temperatures $T_{L}$ and $T_{R}$ with respect to the Hamiltonian $H_{L}$ and $H_{R}$. Thus, we have the canonical states
$\rho_{\beta_{k}}=\exp(-\beta_{k}H_{k})/Z,$ (4)
where $k=L,R$ denotes the subsystem, $\beta_{k}=1/T_{k}$ the inverse temperature and $Z_{k}=\text{tr}[\exp(-\beta_{k}H_{k})]$ the respective partition function. Here, we have set $k_{B}=1$.
We investigate the energy current for various subsystem configurations, including the coupling of (i) two XX chains, (ii) two critical Ising chains, (iii) one XX chain to a critical Ising chain, and (iv) two XXZ chains. We introduce all models in the following.
Figure 2: Time evolution of the energy current for temperatures $T_{L}/J=1/6,T_{R}/J=1/8$, total system size $L=28$ for the coupling of (a) two XX chains, (b) two critical Ising chains and (c) one XX chain to a critical Ising chain. Numerical data for two different pure random initial states (NN=1) and for the average over many initial states (N=100) are compared to the CFT value. Note that larger fluctuations at longer times do not contradict the bound in Eq. (20), since this bound is not tight.
III Models
In our paper, we examine three different examples of spin-1/2 models, which correspond either to noninteracting systems or interacting integrable systems. All of them feature ballistic transport of energy. For simplicity, we focus on the homogeneous case and comment on the subsystem configuration afterwards. All models are one-dimensional lattice models of the type
$H=\sum_{r=1}^{L-1}h_{r}\;,$ (5)
where $h_{r}$ are local Hamiltonians and $L$ denotes again the number of lattice sites.
First, we introduce the spin-$1/2$ XXZ model. For this model, the local Hamiltonians are given by Bertini et al. [2021]
$h_{\text{XXZ},r}=J\left(S_{r}^{x}S_{r+1}^{x}+S_{r}^{y}S_{r+1}^{y}+\Delta S_{r}^{z}S_{r+1}^{z}\right)\;,$ (6)
where $S_{r}^{i}\ (i=x,y,z)$ are the spin-$1/2$ operators at site $r$, $J>0$ is the antiferromagnetic exchange coupling constant, and $\Delta$ is the anisotropy in $z$ direction. Using the Jordan-Wigner transformation Jordan and Wigner [1928], the spin-$1/2$ XXZ chain can be mapped onto a chain of spinless fermions. After this transformation, the local Hamiltonians in Eq. (6) take on the form
$h_{\text{XXZ},r}=$ $\frac{J}{2}\big(c_{r}^{\dagger}c_{r+1}+\text{h.c.}\big)$
$+$ $J\Delta\left(n_{r}-\frac{1}{2}\right)\left(n_{r}-\frac{1}{2}\right)\;,$
where $c_{r}^{\dagger}$ and $c_{r}$ are fermionic creation and annihilation operators at site $r$, and $n_{r}=c_{r}^{\dagger}c_{r}$ denotes the corresponding occupation number operator.
Figure 3: Similar data as the one in Fig. 2 but now for different total system sizes $L$. Here, the numerical data are again averaged over $N=100$ pure random initial states. For larger system sizes $L$, finite size effects are visible for later times tJtJ.
The second model we consider is the special case of the spin-$1/2$ XXZ chain with $\Delta=0$, i.e., the XX chain. The local Hamiltonian reads
$h_{\text{XX},r}=J\left(S_{r}^{x}S_{r+1}^{x}+S_{r}^{y}S_{r+1}^{y}\right)\;.$ (8)
This model is equivalent to a free-fermion chain, as can be seen by setting $\Delta=0$ in Eq. (III) above.
The last spin model we investigate in this paper is the transverse-field Ising model. This model is described by the local Hamiltonian
$h_{\text{Ising},r}=JS_{r}^{x}S_{r+1}^{x}+\frac{h}{2}(S_{r}^{z}+S_{r+1}^{z})\;,$ (9)
where the constant $h$ denotes the strength of the transverse magnetic field. Due to the Jordan-Wigner transformation, the transverse-field Ising model is, up to a constant, equivalent to the Kitaev chain Lieb et al. [1961], Pfeuty [1970], Kitaev [2001], Suzuki et al. [2013],
$h_{\text{Kitaev},r}=\frac{J}{4}\Big(c_{r}^{\dagger}-c_{r}\Big)\Big(c_{r+1}^{\dagger}+c_{r+1}\Big)+hn_{r}\;.$ (10)
Throughout our paper, we focus on the critical transverse-field Ising chain. Therefore, we set $h=1/2$.
As mentioned earlier, we investigate total systems with $L$ sites, composed of two subsystems with $L_{L}=L_{R}=L/2$ sites each. As indicated in Fig. 1, we slightly change notation and use for the left subsystem the sites $r=-L_{L}+1,\dots,0$ and for the right system the sites $r=1,\dots,L_{R}$. The left subsystem interacts with the right subsystem at the sites $r=0$ (left half) and $r=1$ (right half). The interaction is described by the Hamiltonian $H_{C}$, which changes for our different configurations. We use in the case of (i) the coupling of two XX chains
$H_{C}=J_{C}(S_{0}^{x}S_{1}^{x}+S_{0}^{y}S_{1}^{y})\;,$ (11)
for the case of the coupling of (ii) two critical Ising chains and (iii) one XX chain to a critical Ising chain the term
$H_{C}=J_{C}S_{0}^{x}S_{1}^{x}\;,$ (12)
and for the coupling of (iv) two XXZ chains the term
$H_{C}=J_{C}\left(S_{0}^{x}S_{1}^{x}+S_{0}^{y}S_{1}^{y}+\Delta_{C}S_{0}^{z}S_{1}^{z}\right)\;.$ (13)
In our paper, we set $J_{C}=1$ for all configurations. All chosen coupling Hamiltonians ensure that energy can flow between both subsystems. Before we proceed to our numerical results for the energy current in the various subsystem configurations, we introduce our dynamical typicality approach. We also briefly discuss the result obtained within CFT Bernard and Doyon [2012].
IV Methods
IV.1 Dynamical typicality
Figure 4: Steady-state energy current $J_{0}$ versus temperature of the right subsystem $T_{R}$ for the coupling of (a) two XX chains, (b) two critical Ising chains, and (c) one XX chain to a critical Ising chain for the total system size $L=24$ and $T_{L}/J=1/8$. The numerical data are compared to the CFT result (25).
We now turn to a discussion of dynamical quantum typicality (DQT) Jin et al. [2021], Heitmann et al. [2020], which is the key concept used in our numerical approach. We first focus on the homogeneous case of a single Hamiltonian $H$ and a single temperature $T$, and afterwards come to the bipartite case with two $H_{L}$, $H_{R}$ and two $T_{L}$, $T_{R}$.
The basic idea of DQT Gemmer and Mahler [2003], Goldstein et al. [2006], Popescu et al. [2006] is that one can approximate the expectation value of an ensemble density matrix,
$\text{tr}[\rho(t)O]\,,\quad\rho(t)=e^{-iHt}\,\rho\,e^{iHt}\,$, (14)
by an expectation value of just one pure state,
$\langle\psi(t)|O|\psi(t)\rangle\,,\quad|\psi(t)\rangle=e^{-iHt}\,|\psi\rangle\,$, (15)
where $O$ is a local operator. For such a local operator, the moments
$M_{n}=\frac{\text{tr}[O^{n}]}{D}$ (16)
with the Hilbert-space dimension $D$, do not depend on system size $L$. Moreover, we assume that these moments are finite.
If $\rho$ was the microcanonical ensemble and $|\psi\rangle$ an energy eigenstate, this idea would be identical to the well-known eigenstate thermalization hypothesis (ETH) Deutsch [1991], Srednicki [1994], Rigol et al. [2008]. In contrast to ETH, however, DQT is no assumption and relies on a particular construction of the pure state $|\psi\rangle$. Specifically, for a given ensemble density matrix $\rho$, this construction reads
$|\psi\rangle=|\psi(0)\rangle\propto\sqrt{\rho}\,|\Phi\rangle\,$, (17)
which is the action of the square-root of $\rho$ on a so-called Haar-random pure state $|\Phi\rangle$ Bartsch and Gemmer [2009]. As discussed below, this construction particularly enables the reduction of memory in numerical simulations, since pure states (of dimension $D$) instead of density matrices (of dimension $D^{2}$) can be used.
The particular pure state $|\Phi\rangle$ does neither depend on $\rho$ nor $H$. In an arbitrarily chosen orthogonal basis, it is given by
$|\Phi\rangle\propto\sum_{m=1}^{D}c_{m}|m\rangle\,$, (18)
where $D$ is again the Hilbert-space dimension and $c_{m}$ are complex coefficients. Both, the real and imaginary part of these coefficients are drawn at random according to a Gaussian probability distribution with zero mean. For such a $|\Phi\rangle$ and the resulting $|\psi\rangle$, DQT then leads to the approximation Reimann [2007], Bartsch and Gemmer [2009]
$\text{tr}[\rho(t)O]=\frac{\langle\psi(t)|O|\psi(t)\rangle}{\langle\psi|\psi\rangle}+\varepsilon\,$, (19)
where the error $\varepsilon$ depends on the particular realization of the random state $|\Phi\rangle$. However, it possible to make a statistical statement and $\varepsilon$ vanishes on average. More importantly, however, the standard deviation is bounded from above,
$\sigma(\varepsilon)\leq b={\cal O}\left(\frac{1}{\sqrt{D_{\text{eff}}}}\right)\,.$ (20)
Here, $D_{\text{eff}}$ is the so-called effective dimension. In the case of the canonical ensemble $\rho\propto\exp(-\beta H)$, which is most relevant in the context of our work and explained in more detail in Appendix A, the effective dimension is given by the partition function
$D_{\text{eff}}=\text{tr}[\exp(-\beta(H-E_{0}))]$ (21)
with the ground-state energy $E_{0}$. Thus, $D_{\text{eff}}$ counts the number of thermally occupied energy eigenstates. In the high-temperature limit $T\to\infty$, $D_{\text{eff}}=D=2^{L}$ and hence the statistical error $\varepsilon$ decreases exponentially fast with system size $L$. In the zero-temperature limit $T\to 0$, however, $D_{\text{eff}}=1$ and thus the statistical error $\varepsilon$ can be in principle large. Still, one has to take into account that the bound in Eq. (20) is not tight. In fact, at $T=0$, it turns out that the statistical error becomes $\varepsilon=0$. This fact is a kind of trivial consequence for a Haar-random pure state $|\Phi\rangle$: It is also random in the energy eigenbasis and the coefficient $c_{0}$ for the ground state $|0\rangle$ might be small but likely nonzero. At $T=0$, only this coefficient survives and one has $|\psi\rangle=|0\rangle$ Balz et al. [2018]. Note that this fact is important, since we have $D_{\text{eff}}$ close to 1 for the low temperatures considered later, in contrast to the large $D_{\text{eff}}$ in usual high-temperature consideratios, e.g., $D_{\text{eff}}=D=2^{30}$ for $L=30$.
Let us next turn to the specific setup in Fig. 1, which is studied in our work. For this setup, we could proceed completely analogous and choose for the construction of the pure state $|\psi\rangle$ in Eq. (17) an ensemble density matrix of the form
$\rho\propto\exp(-\beta_{L}H_{L})\otimes\exp(-\beta_{R}H_{R})\,.$ (22)
While this choice is valid, the resulting $|\psi\rangle$ would be no product state of the two halves of the system, because $|\Phi\rangle$ is drawn at random from the full Hilbert space and thus highly entangled. Hence, to implement the product structure, we proceed slightly different and choose the pure state as
$|\psi\rangle\propto\exp\left(-\frac{\beta_{L}H_{L}}{2}\right)|\Phi_{L}\rangle\otimes\exp\left(-\frac{\beta_{R}H_{R}}{2}\right)|\Phi_{R}\rangle\,$, (23)
where $|\Phi_{L}\rangle$ and $|\Phi_{R}\rangle$ are now drawn at random from the Hilbert space of the left and right half. While this choice does not change the typicality relation in Eq. (24), it goes along with a larger statistical error $\varepsilon$, which is given by the partition functions of the two halves, rather than the partition function of the entire system. To further reduce the statistical error, we average over $N=100$ different realizations of $|\psi\rangle$,
$\text{tr}[\rho(t)O]=\frac{1}{N}\sum_{i=1}^{N}\frac{\langle\psi_{i}(t)|O|\psi_{i}(t)\rangle}{\langle\psi_{i}|\psi_{i}\rangle}\,$, (24)
if not stated otherwise. In this way, the statistical error in Eq. (20) is multiplied by the factor $1/\sqrt{L}$ Jaklič and Prelovšek [2000]. In our case, the local observable $O$ is mostly the contact current $J_{0}$, but we also consider currents $J_{r}$ at other sites $r$ as well as the local energy densities $h_{r}$.
In our numerical simulations, we calculate the action of the imaginary-time operators $e^{-\beta_{L}H_{L}/2}$ and $e^{-\beta_{R}H_{R}/2}$ on the Haar-random pure states $|\Phi_{L}\rangle$ and $|\Phi_{R}\rangle$ with a fourth-order Runge-Kutta scheme, where we choose a very small step $\delta\beta J=0.0001$. This scheme is also used for the action of the real-time operator $e^{-iHt}$ on the pure state $|\psi\rangle$, where we choose a larger but still small step $\delta tJ=0.01$. By doing so and also using sparse-matrix techniques, we can treat system sizes outside the range of standard exact diagonalization. Alternatively, other propagation methods, such as Trotter decompositions Suzuki [1985], De Raedt [1987], Berry et al. [2006] or Chebyshev polynomial expansions Tal-Ezer and Kosloff [1984], Weiße et al. [2006], can be used as well.
IV.2 Conformal field theory
Figure 5: Steady-state profile for (a) the local density $\langle h_{r}\rangle$ and (b) the local current $\langle J_{r}\rangle$ for the coupling of two XXZ chains with the same anisotropies $\Delta$, $T_{L}/J=1$$T_{R}/J=1/2$, and total system size $L=24$. (Note that in the $x$-axis the position $r$ is divided by the time $t$.) The numerical data are compared to the results from GHD in Ref. Bertini et al. [2016]. Here, the numerical data are shown for both, a random product state (open symbols) and a random state in the entire Hilbert space (closed symbols).
The energy transport in a bipartite CFT setup has been studied by Bernard and Doyon Bernard and Doyon [2012]. Specifically they considered a CFT separated into two halves prepared at different temperatures. These halves were then glued together and the non-equilibrium steady state emerging at late times was studied. In particular, for the energy current they obtained the universal formula
$J_{E}=\frac{\pi c}{12}(T_{L}^{2}-T_{R}^{2})\;,$ (25)
where $c$ denotes the central charge of the CFT. Given that CFTs generically describe the universal low-energy behavior of certain lattice models, the result (25) is expected to hold only in the low-temperature regime. The central charge $c$ is a specific characteristic of the CFT Di Francesco et al. [1997], physically it can be interpreted as the number of degrees of freedom, which here contribute to the energy transport through the system.
The setup leading to the universal CFT result (25) possesses translation invariance after the gluing quench. As an extension, Ref. Fischer et al. [2020] studied the coupling of an XX chain to a critical Ising chain, i.e., our setup (iii) above. Using both the non-equilibrium Green function formalism and the density matrix renormalization group method, the energy current was found to behave as
$J_{E}=\frac{\pi}{12}\text{min}(c_{L},c_{R})\,(T_{L}^{2}-T_{R}^{2})\;.$ (26)
This result can be viewed as a bottleneck effect, where the subsystem with the lower number of degrees of freedom as reflected by the lower central charge limits the energy transport. The result (26) is also consistent with a purely field theoretic analysis by Bernard et al. Bernard et al. [2015], where, however, the precise link to a microscopic setup remains elusive.
V Results
Figure 6: Similar data as the one in Fig. 4 but now for the coupling of two XXZ chains with anisotropies (a) $\Delta=1$, (b) $\Delta=0.5$, and (c) $\Delta=0.25$.
Now, we turn to our numerical results, where we compare the energy current for various subsystem configurations to the value from CFT in Eq. (25). We first focus on the coupling of (i) two XX chains, (ii) two critical Ising chains, and (iii) one XX chain to a critical Ising chain. In CFT, the XX chain has the central charge $c=1$ while the transverse-field Ising chain at the critical point has the central charge $c=1/2$ Di Francesco et al. [1997].
To verify our numerical method initially, we compute the time evolution of the energy current $J_{0}$ for our considered subsystem configurations for systems with $L=28$ sites and a fixed temperature of the left ($T_{L}/J=1/6$) and right ($T_{R}/J=1/8$) subsystem in Fig. 2. All subsystem configurations exhibit a transition regime characterized by oscillations and a buildup of correlations across the interface. After a specific timescale, the current reaches an approximately constant plateau, indicating the establishment of a non-equilibrium steady state. Also in Fig. 2, we use two different realizations of a single random pure state and compare them with the average over $N=100$ random pure states. With the averaging, the statistical error reduces remarkably, especially for the long-term behavior of the energy current $J_{0}$. Furthermore, the steady state of the energy current agrees nicely with the value from CFT for all subsystem configurations. In the following, we average all our numerical results over $N=100$ random pure initial states, if not stated differently.
Next, we investigate the dependence of the energy current on the total system size. We thus depict the time evolution of the energy current $J_{0}$ for varying total system sizes $L$ and for the fixed initial temperatures of $T_{L}/J=1/6$ and $T_{R}/J=1/8$ in Fig. 3. For longer times, finite-size effects are visible for smaller $L$. Nevertheless, below the time where finite sizes effects are relevant, the data are close to the steady-state value according to CFT, especially for the coupling of an XX chain to the critical Ising chain, cf. Fig. 3 (c). In the following, we focus on total system sizes of $L=24$, where finite-size effects and deviations from the CFT value are sufficiently small.
As one of our central results, we now calculate the steady-state energy current $J_{0}$ for a fixed initial temperature of $T_{L}/J=1/8$ of the left subsystem and different initial temperatures of the right subsystem, see Fig. 4. We compare our numerics to the CFT values in Eq. (25) for $c=1/2$ and $c=1$. For all considered subsystem configurations (i) – (iii), we have a convincing agreement between our numerical results and CFT. Also our data do not significantly depend on system size, as we show in Fig. 4 by also calculating the energy current for exemplary temperatures $T_{R}/J$ of the right subsystem for a total system size of $L=26$. Furthermore, for the coupling of an XX chain to a critical Ising chain, the energy current is determined by the system with the lower central charge, which is, in this scenario, the critical Ising chain, see Fig. 4 (c). Consequently, we observe a bottleneck effect Fischer et al. [2020], where the energy transport is limited by the subsystem with the fewer degrees of freedom.
Let us also give a technical remark here. It would be certainly interesting to explore even lower temperatures, e.g., to detect a region, where eventually DQT and CFT differ significantly from each other. However, in Fig. 4 we already consider a temperature $T_{R}/J$ = 0.05 or inverse temperature $\beta_{R}J=20$. To do so, we have to accurately approximate a small number of order $e^{-20}$ by an iterative procedure in imaginary time. An accurate approximation for even lower temperatures is quickly beyond the scope of such an interative procedure.
As the final subsystem configuration, we proceed to the (iv) coupling of two XXZ chains. We first focus on the case where all anisotropies are equal, i.e., $\Delta=\Delta_{L}=\Delta_{R}=\Delta_{C}$. We again benchmark our numerical calculation against existing results in the literature. In Fig. 5, we compare our numerics for a total system with $L=24$ sites, fixed temperatures of the left and right subsystem, and different anisotropies $\Delta$ to results from GHD. The numerical data for the site dependence of the local energy density (Fig. 5 (a)) as well as the data for the site dependence of the local energy current (Fig. 5 (b)) matches the results of Ref. Bertini et al. [2016] convincingly. Here, the numerical data are shown for both, a random product state (open symbols) and a random state in the entire Hilbert space (closed symbols). As shown in Fig. 5, the statistical error for a random state in the entire Hilbert space is slightly smaller than the statistical error for a random product state. It would be also interesting to analyze states, which interpolate between these two states of no and maximum bipartite entanglement. But we leave this analysis for future research.
Focusing again on the contact site, we calculate the energy current $J_{0}$ for the coupling of two XXZ chains with total system size $L=24$, fixed temperature $T_{L}/J=1/8$, and the anisotropies $\Delta=0.25,0.5$, and $1.0$, see Fig. 6. For $1\leq\Delta<1$, the ground state of the XXZ chain is gapless. The low-energy behavior of the XXZ chain in this regime is described by a CFT with central charge of $c=1$ Giamarchi [2003]. Again, we recover the temperature dependence in Eq. (25) for all considered anisotropies, and our data agrees with the CFT result convincingly and also do not significantly depend on system size.
At last, we also calculate the energy current $J_{0}$ for the coupling of two XXZ chains, but now with the anisotropies $\Delta_{L}=0$ and $\Delta_{R}=\Delta_{C}=0.5$. Additionally, in this case, we find convincing agreement with the CFT value, as shown in Fig. 7. Overall, this and the previous results demonstrate that our typicality approach is suitable to capture the energy current down to low temperatures.
Figure 7: Similar data as the one in Fig. 6 but now for the coupling of two XXZ chains with anisotropies $\Delta_{L}=0$ and $\Delta_{R}=\Delta_{C}=0.5$.
VI Conclusion
To summarize, we discussed the method of dynamical quantum typicality, mostly used in linear response theory and at high temperatures, in less explored contexts. First, we applied it to energy flow between systems composed of two spin-chain domains at different temperatures and, second, to low temperatures. To test our method, we investigated the energy current in different spin-$1/2$ subsystem configurations, namely the coupling of (i) two XX chains, (ii) two critical Ising chains, (iii) one XX chain to a critical Ising chain, and (iv) two XXZ chains. We compared our numerical results to results from CFT and GHD. For all considered subsystem configurations and temperatures, we found a convincing agreement between our quantum typicality approach and these established methods. Therefore, we demonstrated the suitability of our approach for the study of bipartite systems at low temperatures.
Interesting directions of future investigations include additional models, e.g., the three-state quantum Potts chain, which in CFT has a central charge of $c=4/5$. In addition, it is desirable to extend the investigation to gapped systems. Our expectation is that these systems are harder to treat by means of DQT, compared to the critial ones studied here.
Acknowledgments
This work was funded by the Deutsche Forschungsgemeinschaft (DFG) under Grant No. 397067869 within the DFG Research Unit FOR 2692 under Grant No. 355031190.
DATA AVAILABILITY
The data that support the findings of this article are openly available data .
Appendix A Details on the error bound
In the following, we provide details on the error bound for the variance, which can be found in, e.g., Knipschild [2021] in similar form.
Let us consider the variance
$\sigma^{2}=\overline{\langle\psi|O|\psi\rangle^{2}}-(\overline{\langle\psi|O|\psi\rangle})^{2}\;,$ (27)
where the overbar denotes the ensemble average over pure states $|\psi\rangle$. For a pure state $|\psi\rangle$ according to Eq. (17) one obtaines the variance
$\sigma^{2}$ $=$ $\sum_{i,j,k,l}\overline{c_{i}^{\ast}c_{j}c_{k}^{\ast}c_{l}}\langle m_{i}|\sqrt{\rho}O\sqrt{\rho}|m_{j}\rangle\langle m_{k}|\sqrt{\rho}O\sqrt{\rho}|m_{l}\rangle$ (28)
$-\sum_{i,k}\langle m_{i}|\sqrt{\rho}O\sqrt{\rho}|m_{i}\rangle\langle m_{k}|\sqrt{\rho}O\sqrt{\rho}|m_{k}\rangle\;.$
We now use the identity $\overline{c_{i}^{\ast}c_{j}c_{k}^{\ast}c_{l}}\approx\delta_{ij}\delta_{kl}+\delta_{il}\delta_{jk}$. Here, we can neglect the case $i=j=k=l$ for large dimensions $D\gg 1$ because this single case has only a contribution of order $\mathcal{O}(1/D)$. With this identity, we can simplify Eq. (28) to
$\sigma^{2}$ $\approx$ $\sum_{i,j}\langle m_{i}|\sqrt{\rho}O\sqrt{\rho}|m_{j}\rangle\langle m_{j}|\sqrt{\rho}O\sqrt{\rho}|m_{i}\rangle\;$ (29)
$=$ $\text{tr}[\rho O\rho O]\,.$
Next, we want to bound $\text{tr}[\rho O\rho O]$ from above. To this end, we use the spectral decomposition $p=\sum_{i}p_{i}|i\rangle\langle i|$ and get
$\text{tr}[\rho O\rho O]=\sum_{i,j}p_{i}p_{j}|\langle i|O|j\rangle|^{2}\;.$ (30)
Using
$(p_{i}-p_{j})^{2}=p_{i}^{2}-2p_{i}p_{j}+p_{j}^{2}\geq 0,$ (31)
we then obtain the inequality
$\text{tr}[\rho O\rho O]$ $\leq$ $\sum_{i,j}\frac{p_{i}^{2}+p_{j}^{2}}{2}|\langle i|O|j\rangle|^{2}$ (32)
$=$ $\sum_{i}p_{i}^{2}\langle i|O^{2}|i\rangle$
$=$ $\text{tr}[\rho^{2}O^{2}]\;.$
In the eigenbasis $|O_{j}\rangle$ of the operator $O$ we obtain for the trace on the r.h.s the estimation
$\text{tr}[\rho^{2}O^{2}]$ $=$ $\sum_{j}\langle O_{j}|\rho^{2}|O_{j}\rangle\langle O_{j}|O^{2}|O_{j}\rangle$ (33)
$\leq$ $O_{\text{max}}^{2}\,\text{tr}[\rho^{2}]\;,$
where $O_{\text{max}}=\text{max}[O_{j}]$ is the maximal eigenvalue. Thus, in summary, the variance can be bounded from above as
$\sigma^{2}\leq O_{\text{max}}^{2}\,\text{tr}[\rho^{2}]\,$, (34)
where $O_{\text{max}}$ of the local operator $O$ does not depend on system size $L$. Inserting for the density matrix $\rho$ the canonical ensemble leads to
$\frac{\sigma^{2}}{O_{\text{max}}^{2}}\leq\frac{\text{tr}[(e^{-\beta H})^{2}]}{(\text{tr}[e^{-\beta H}])^{2}}$ (35)
or, equivalently,
$\frac{\sigma^{2}}{O_{\text{max}}^{2}}\leq\frac{\text{tr}[(e^{-\beta(H-E_{0})})^{2}]}{(\text{tr}[e^{-\beta(H-E_{0})}])^{2}}=\frac{\text{tr}[e^{-2\beta(H-E_{0})}]}{(\text{tr}[e^{-\beta(H-E_{0})}])^{2}}$ (36)
with the ground-state energy $E_{0}$. Because
$\text{tr}[e^{-2\beta(H-E_{0})}]\leq\text{tr}[e^{-\beta(H-E_{0})}]\,$, (37)
one gets the expression
$\frac{\sigma^{2}}{O_{\text{max}}^{2}}\leq\frac{\text{tr}[e^{-\beta(H-E_{0})}]}{(\text{tr}[e^{-\beta(H-E_{0})}])^{2}}=\frac{1}{\text{tr}[e^{-\beta(H-E_{0})}]}\,.$ (38)
By introducing the effective dimension
$D_{\text{eff}}=\text{tr}[e^{-\beta(H-E_{0})}]\,$, (39)
this expression can be written as
$\sigma^{2}\leq O_{\text{max}}^{2}\,\frac{1}{D_{\text{eff}}}\,.$ (40)
Omitting the $L$-independent $O_{\text{max}}$ finally yields
$\sigma\leq b={\cal O}\left(\frac{1}{\sqrt{D_{\text{eff}}}}\right)\,$, (41)
cf., e.g., Jaklič and Prelovšek [2000]. Note that this bound can be derived analogously for a time-dependent operator $O(t)$ in the Heisenberg picture, because (i) $O(t)$ remains Hermitian under time evolution and (ii) the spectrum and thus $O_{\text{max}}$ is invariant under time evolution. Therefore, the bound in inquality (34) holds for $O(t)$ with the same constant $O_{\text{max}}$.